A fishing boat travels at 20 mph on its way to the fishing grounds. It takes two hours to reach the grounds. On the way back, when the boat is loaded with fish, it can only travel at 15 mph. How long will it take the boat to get home?

Answer by checkley79(3341)

D=RT
D=20*2
D=15X
20*2=15X
40=15X
X=40/15
X=2.667 HOURS OR 2 HOURS 40 MINUTES.